1)How many liters of pure acid must be added to 40 L of a 20% acid solution to make a solution which is 36% acid?
2)A laboratory technician has a solution that is 25% acid. He wishes to increase the concentration to 40% by adding pure acid. If he has 800 mL of solution to begin with, how much acid must be added to make the desired concentration?

Answer by checkley77(12844)

x+.2*40=.36(40+x)
x+8=14.4+.36x
x-.36x=14.4-8
.64x=6.4
x=6.4/.64
x=10 liters of pur acid is needed.
Proof:
10+.2*40=.36(40+10)
10+8=.36*50
18=18
-----------------------------------------------
.25*800+x=.4(800+x)
200+x=320+.4x
x-.4x=320-200
.6x=120
x=120/.6
x=200 ml of pure acid is needed.
Proof:
.25*800+200=.4(800+200)
200+200=.4(1,000)
400=400